A man is three times older than his son. If the differences between the two of them is 20, what is the sum total of the age?
Answer by suruman(21)

Let age of the son be x
Since man is 3 times as old as son, man's age = 3x
Difference between their age = 20.
Thus,
3x - x = 20
2x = 20
x = 10
Thus age of son = 10
Age of the man = 3*10 = 30
Thus sum of their ages = 30 + 10 = 40